OpenStudy (anonymous):

Multiply and simplify. 4c^2/4c^2-32c+65*4c-16/2c please help me by showing the steps on how to do this. Thanks!

OpenStudy (anonymous):

can you use some parentheses to clarify the question

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

\[(4c^2/4c^2-32c+64)\times(4c-16/2c)\]

OpenStudy (anonymous):

\[\frac{4c^2}{4c^2-32c+64}*\frac{4c-16}{2c}\]

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

I'd first factor a 4 out of the 1st fraction

OpenStudy (anonymous):

So 1/-32c+64 ?

OpenStudy (anonymous):

Now the 1st denominator is easy to factor

OpenStudy (anonymous):

Oh I see. Let me see how far I get doing that.

OpenStudy (anonymous):

i'm sorry actually the 4's would cancel, so you don't multiply the entire thing by 4

OpenStudy (anonymous):

\[\frac{4c^2}{4(c^2-8c+16)}*\frac{4(c-4)}{2c}\]

OpenStudy (anonymous):

can you work on it from there?

OpenStudy (anonymous):

Let me try :)

OpenStudy (anonymous):

Do I multiply straight across? 4c^2*4(c-4)?

OpenStudy (anonymous):

not yet, factor the denominator of the 1st fraction first, you'll see why you shouldn't multiply yet

OpenStudy (anonymous):

Oh! As if it is a quadratic?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4(c-4)^2 ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now you can start cancelling stuff

OpenStudy (anonymous):

\[\frac{4c^2}{4(c^2-8c+16)}*\frac{4(c-4)}{2c}=\frac{\cancel{4}c^2}{\cancel{4}\cancel{(c-4)}(c-4)}*\frac{4\cancel{(c-4)}}{2c}\]

OpenStudy (anonymous):

this can be further simplified

OpenStudy (anonymous):

now you multiply across

OpenStudy (anonymous):

you there moonie?

OpenStudy (anonymous):

Yeah I'm trying to do the math. :) 4c^2/2c^2-4 ?

OpenStudy (anonymous):

At this point we have this: \[\frac{4c^2}{2c(c-4)}\]

OpenStudy (anonymous):

Rob, thank you but the original equation was written incorrectly, Rickjbr is heping me get the original equation solved correctly.

OpenStudy (anonymous):

Oh I see, I can't square the c's.

OpenStudy (anonymous):

you can, but you would have to undo it for the next step, so it's pointless. Also, 2c(c-4)=2c^2-8c

OpenStudy (anonymous):

how did you get 2c^2-4?

OpenStudy (anonymous):

Bad math... I was trying to use the distributive law but I didn't follow all the way through with it.

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

You always want to cancel out as many terms as you can before multiply anything.

OpenStudy (anonymous):

Can you reduce the fraction we have at this point?

OpenStudy (anonymous):

I think so...

OpenStudy (anonymous):

2/-8c ?

OpenStudy (anonymous):

no, walk me through how you got that

OpenStudy (anonymous):

Be cause 2c^2 will go into 4c^2 one time with a remainder of 2? (God I'm horrible at this)

OpenStudy (anonymous):

This is why i told you not to expand the fraction yet, because you can't reduce fractions with addition or subtraction signs in the middle

OpenStudy (anonymous):

\[\frac{4c^2}{2c(c-4)}\]

OpenStudy (anonymous):

This is what we're working with right now

OpenStudy (anonymous):

OH!..... Okay let me try again.

OpenStudy (anonymous):

Are you using this programs eqiation editor? (Just a side question)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So would I expand the numerator since it is squared and then try to cancel from there?

OpenStudy (anonymous):

the numerator can't be expanded

OpenStudy (anonymous):

What is 4/2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

So coefficients first then variables and exponents?

OpenStudy (anonymous):

Yes, but everything has to be in factored form for you to divide and cancel. It is in factored form right now, so you can do that.

OpenStudy (anonymous):

rob, you should read the 5th post from the top :/

OpenStudy (anonymous):

making any progress?

OpenStudy (anonymous):

Um, maybe... 4c^2/2c(c-4) = 2c^2/c(c-4) Can I use the distributive law on the denominator now?

OpenStudy (anonymous):

nope, there's some more cancelling to d o

OpenStudy (anonymous):

Is my equation correct so far?

OpenStudy (anonymous):

yes, it is equivalent

OpenStudy (anonymous):

What is \[\frac{c^2}{c}\]

OpenStudy (anonymous):

Do I divide 2 by -4?

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

no you can't do that

OpenStudy (anonymous):

Properties of exponents, quotient rule: \[\huge \frac{a^m}{a^n}=a^{m-n}\]

OpenStudy (anonymous):

\[\huge \frac{c^2}{c^{1}}=c^{2-1}=c^1=c\]

OpenStudy (anonymous):

Oh, I didn't know I could use 1 as an exponent

OpenStudy (anonymous):

If that is the case then we are left with 2c/ (c-4) ?

OpenStudy (anonymous):

You can't touch anything inside of the parentheses because there are addition and subtraction symbols in there

OpenStudy (anonymous):

Yes, that's the final answer, no further factoring possible

OpenStudy (anonymous):

Omg that was worse then pulling a tooth. Thank you so much for your patience!

OpenStudy (anonymous):

you're welcome.

OpenStudy (anonymous):

1 is an implied exponent, nobody writes it because it's always there

OpenStudy (anonymous):

I understand. I just don't remember things like that. Okay well, I have four more problems to get figured out. Thank you again!

OpenStudy (anonymous):

Refer to the solution attachment.